Determine the maximums and minimums of the following function: x2 + 2y2 - xy
Answer #1

fix,y) = 2 + 2y²-ny tet -Diff ear worotn 20-y 2 ny 2n-y = 0 =) Y=24 Je ne 4 4 4 g2f = 4 = l = 2 af = 4y - 1 = 24-120 = 4(2n)-